What is the bond prices?

What is the zero discount rate?

What is the forward rate?

Par Value=1000

Maturity (yrs) YTM Coupon Rate
0.5 0.06 0.062
1.0 0.07 0.072
1.5 0.08 0.082
2 0.09 0.092

The market convention is semi-annual coupons and semi-annual compounding
Answer #1

Ans: 1) Value of Bond

Maturity (yrs) YTM Coupon Rate Interest Present Value Factor Present value
0.5 0.06 0.062 $62.00 0.97 $60.19
1 0.07 0.072 $72.00 0.94 $67.54
1.5 0.08 0.082 $82.00 0.88 $71.81
2 0.09 0.092 $92.00 0.76 $70.21
$269.75

Present value factor = ((1/(1+R1/2))^1)*((1/(1+R2/2))^2)*((1/(1+R3/2))^n)

Here R1 is Coupon rate of first six month and R2 is Interest of next six month

Computation of bond price = Issue price of the bond /( 1+ Yield till maturity) + Sum of (Interest received / ( 1+ yield till maturity))for N number of period

Computation of bond price = Present value of the bond + Present value of Interest received

Present value of Principle amount to be received = Redeemable value of the bond * Present value factor

= $1,000*0.76

= $760

Value of Bond = $760 + $269.75

= $1029.75

2) Zero discount rate

Present value of zero discounted Bond = Redeemable value of the bond * Present value factor

   1,000*(1+(1+r/2))^4 = $1,000*0.76

(1+(1+r/2))^4 = $760/1,000

  (1+(1+14%/2))^4= 0.76

So Zero Discounted rate = 14%

3) Forward rate

   Present value of Bond= Face value of the bond / ( (1+ Interest rate)*(1+Forward rate))

Maturity (yrs) YTM/Forward rate Coupon Rate
0.5 0.6

0.62

1

0.7

0.72
1.5 0.8 0.82
2 0.9 0.92